Please help me solve this problem..I will appreciate your help.
The random variable X has the following distribution.

 X         p(X)
---------------
-1         3c
 0         2c
 1         0.4
 2         0.1

Find the value of constant c.

Found 2 solutions by Edwin McCravy, natolino_2017:
Answer by Edwin McCravy(20062)

The random variable X has the following distribution.

 X         p(X)
---------------
-1         3c
 0         2c
 1         0.4
 2         0.1

Find the value of constant c.

The sum of the probabilities must = 1, so the equation is

3c + 2c + 0.4 + 0.1 = 1
           5c + 0.5 = 1
                 5c = 0.5
                  c = 0.1

That's the answer.  That means the distribution is

 X         p(X)
---------------
-1         0.3
 0         0.2
 1         0.4
 2         0.1
---------------
     sum = 1.0

Edwin

Answer by natolino_2017(77)

The sum of every proability must be equal to one.
So 3c + 2c + 0.4 +0.1 = 1
5c = 0.5
c=0.1
@natolino_